OpenStudy (anonymous):

Which pair of ratios can form a true proportion? A 9/12 , 5/8 B 25/20 , 16/21 C 6/27 , 2/9 D 7/12 , 10/15

OpenStudy (texaschic101):

proportions are nothing but equivalent fractions so reduce all the fractions and see which ones are equal 9/12 = 3/4 5/8 does not reduce....this is not a proportion 25/20 = 5/4 16/21 does not reduce....this is not a proportion 6/27 = 2/9 2/9 = 2/9 <==== BINGO...these are proportional 7/12 -- this does not reduce 10/15 = 3/5......this is not a proportion